OpenStudy (anonymous):

Express this equation in logarithmic form. 6^3 = 216

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Oh okay so here it's important to remember the rule \[\large ln(a)^b = b \times ln(a)\] so for you here \[\large 6^3 = 216\] will then become (after taking the log of both sides) \[\large 3 \times \ln(6) = \ln(216)\] Or \[\large 3 = \frac{\ln(216)}{\ln(6)}\]